# Attention Is All you Need For Blind Room Volume Estimation

### v1

@@ Abstract
In recent years, dynamic parameterization of acoustic environments has raised increasing attention in the field of audio processing. One of the key parameters that characterize the local room acoustics in isolation from orientation and directivity of sources and receivers is the geometric room volume. Convolutional neural networks (CNNs) have been widely selected as the main models for conducting blind room acoustic parameter estimation, which aims to learn a direct mapping from audio spectrograms to corresponding labels. With the recent trend of self-attention mechanisms, this paper introduces a purely attention-based model to blindly estimate room volumes based on single-channel noisy speech signals. We demonstrate the feasibility of eliminating the reliance on CNN for this task and the proposed Transformer architecture takes Gammatone magnitude spectral coefficients and phase spectrograms as inputs. To enhance the model performance given the task-specific dataset, cross-modality transfer learning is also applied. Experimental results demonstrate that the proposed model outperforms traditional CNN models across a wide range of real-world acoustics spaces, especially with the help of the dedicated pretraining and data augmentation schemes. ††This work was supported by the National Natural Science Foundation of China under Grant No. 61971015 and Beijing Natural Science Foundation (No.L233032, L223033).
@@ 1 Introduction
Dynamic parameterization of acoustic environments that users evolve has become an emerging topic in recent years. Parameters that characterize local rooms or other acoustic spaces can be used to model or design audio filters for various applications, e.g. speech dereverberation for automatic speech recognition (ASR) and voice communication [1, 2], spatial sound systems with room equalization [3, 4] and etc. In particular, for the proper realization of audio augmented reality (AAR), virtual acoustic objects are required to be seamlessly integrated into the real environment, which makes a good match between acoustical properties of virtual elements and the local space a necessity [5].
@@
where denotes the total area of the room’s surfaces and is the area-weighted mean absorption coefficient in octave band . In practice, in-situ measurements of RIRs and volumes of users’ local acoustic spaces are typically difficult to be carried out. Alternatively, an attractive option is to blindly estimate room acoustic parameters from audio recordings using microphones. The 2015 ACE challenge [7] sets the bar for the blind estimation of RT and DRR from noisy speech sequences. Meanwhile, room volume estimation has long been formulated as a classification problem [8, 9]. With recent advancements in DNNs, formulating the blind room volume estimation as a regression problem by taking advantage of convolutional neural network (CNN) models in conjunction with time-frequency representations has become increasingly relevant. Genovese et al. [10] deployed a CNN trained using both simulated as well as real RIRs, and results show that it can estimate a broad range of volumes within a factor of 2 on real-measured data from the ACE challenge [7]. Similar CNN-based systems were proposed to blindly estimate room acoustic parameters from single-channel [11, 12, 13] or multi-channel [14] speech signals and demonstrated promising results in terms of both estimation accuracy and robustness to temporal variations in dynamic acoustic environments. In addition to the log-energy calculation of spectro-temporal features that prior works generally relied on, Ick et al. [15] introduced a series of phase-related features and demonstrated clear improvements in the context of reverberation fingerprint estimation on unseen real-world rooms.
CNNs are widely considered in the fore-mentioned approaches due to their suitability for learning two-dimensional time-frequency signal patterns for end-to-end modelling. In order to better capture long-range global context, CNN-attention hybrid models that concatenate CNN with a self-attention mechanism have achieved cutting-edge results for numerous tasks such as acoustic event classification [16, 17] and other audio pattern recognition topics [18, 19]. Gong et al. [20] took one step further and devised purely attention-based models for audio classification. The devised Audio Spectrogram Transformer (AST) was assessed on various audio classification benchmarks with new state-of-the-art results, which shows that CNNs are not indispensable in this context.
@@ 2.1 Featurization
Similar to prior literature [10, 11], the Gammatone ERB filterbank is selected for generating time-frequency representation as it leads to low model complexity while preserving signal information that is relevant to this problem. Specifically, the Gammatone filterbank consists of 20 bands covering the range from 50Hz to 2000Hz. We compute the STFT of the audio using a Hann window of 64 samples and a hop size of 32 samples, followed by convolution with the filterbank. This convolution generates a spectral feature (). Additionally, the phase information obtained from the STFT is retained. The phase angles computed for each time-frequency bin can be used to generate a phase feature, and it is then truncated to only include the bands corresponding to frequencies below 500 Hz (i.e. ) as lower frequency behavior generally carries more information corresponding to the room volume [6]. Furthermore, the first-order derivative of phase coefficients along the frequency axis is also concatenated (i.e. ). The configuration of this feature set aligns with the “+Phase” model outlined in [15], which is shown to outperform other methods that rely solely on magnitude-based spectral features.
@@ 4.1 Experimental Design
During the model training phase, MSE was used as the loss function and the Adam optimizer from PyTorch was deployed for optimization. 80 epochs were run in each training session for all models as good convergence behaviors can already be observed for both training and validation. L2 regularization was applied to mitigate potential over-fitting. Additionally, an adaptive learning rate strategy was adopted to ensure efficient convergence during training. For comparison test purposes, we switched between Dataset I and Dataset II, as well as whether or not to use the pretrained model from ImageNet, while maintaining same hyperparameters between the attention-based and CNN-based model to ensure uniformity in model configurations.
@@ 4.2.1 Base Systems
It can be clearly seen that the proposed model outperforms the CNN model in terms of prediction accuracy, relationship with ground truth values and predictive capability. This indicates that neural networks purely based on attention are sufficient (or even more superior than CNN models) to accurately learn the relationship between indoor acoustic characteristics and room volumes, even with the low-layer network configuration and a relatively small number of training epochs.
@@ 4.2.2 Enhanced Systems With Pretraining
With Dataset I, the deployment of the ImageNet pretraining elevated the proposed method’s performance to a new level, yielding a significantly improved room volume estimation accuracy. With the application of the SpecAugment method, Dataset II facilitated to further enhance system performance for all three models. Particularly, the augmentation effect was more prominent in the proposed method w/ Pretrain, confirming the effectiveness of SpecAugment in terms of alleviating overfitting and enhancing models’ generalizability. As a more illustrative example, the best-performing system, i.e. “Proposed method w/ Pretrain” model with Dataset II, resulted in a median and mean absolute error of 223 and 1501 in linear scale respectively, given that the range of test set room volumes was [12, 21000] . In contrast, the median and mean absolute error of the CNN-based system with Dataset II was 524 and 2400 , respectively.

### v3

@@ Abstract
In recent years, dynamic parameterization of acoustic environments has raised increasing attention in the field of audio processing. One of the key parameters that characterize the local room acoustics in isolation from orientation and directivity of sources and receivers is the geometric room volume. Convolutional neural networks (CNNs) have been widely selected as the main models for conducting blind room acoustic parameter estimation, which aims to learn a direct mapping from audio spectrograms to corresponding labels. With the recent trend of self-attention mechanisms, this paper introduces a purely attention-based model to blindly estimate room volumes based on single-channel noisy speech signals. We demonstrate the feasibility of eliminating the reliance on CNNs for this task and the proposed Transformer architecture takes Gammatone magnitude spectral coefficients and phase spectrograms as inputs. To enhance the model performance given the task-specific dataset, cross-modality transfer learning is also applied. Experimental results demonstrate that the proposed model outperforms traditional CNN models across a wide range of real-world acoustics spaces, especially with the help of a dedicated pretraining and data augmentation schemes. ††This work was supported by the National Natural Science Foundation of China under Grant No. 61971015 and Beijing Natural Science Foundation (No.L233032, L223033).
@@ 1 Introduction
Dynamic parameterization of acoustic environments that users evolve has become an emerging topic in recent years. Parameters that characterize local rooms or other acoustic spaces can be used to model or design audio filters for various applications, e.g. speech dereverberation for automatic speech recognition (ASR) and voice communication [1, 2], spatial sound systems with room equalization [3, 4] etc. In particular, for the proper realization of audio augmented reality (AAR), virtual acoustic objects are required to be seamlessly integrated into the real environment, which makes a good match between acoustical properties of virtual elements and the local space a necessity [5].
@@
where denotes the total area of the room’s surfaces and is the area-weighted mean absorption coefficient in octave band . In practice, in-situ measurements of RIRs and volumes of users’ local acoustic spaces are typically difficult to be carried out [7]. Alternatively, an attractive option is to blindly estimate room acoustic parameters from audio recordings using microphones. The 2015 ACE challenge [8] sets the bar for the blind estimation of RT and DRR from noisy speech sequences. Meanwhile, room volume estimation has long been formulated as a classification problem [9, 10]. With recent advancements in DNNs, formulating the blind room volume estimation as a regression problem by taking advantage of convolutional neural network (CNN) models in conjunction with time-frequency representations has become increasingly relevant. Genovese et al. [11] deployed a CNN trained using both simulated as well as real RIRs, and results show that it can estimate a broad range of volumes within a factor of 2 on real-measured data from the ACE challenge [8]. Similar CNN-based systems were proposed to blindly estimate room acoustic parameters from single-channel [12, 13, 14] or multi-channel [15] speech signals and demonstrated promising results in terms of both estimation accuracy and robustness to temporal variations in dynamic acoustic environments. In addition to the log-energy calculation of spectro-temporal features that prior works generally relied on, Ick et al. [16] introduced a series of phase-related features and demonstrated clear improvements in the context of reverberation fingerprint estimation on unseen real-world rooms.
CNNs are widely considered in the aforementioned approaches due to their suitability for learning two-dimensional time-frequency signal patterns for end-to-end modelling. In order to better capture long-range global context, CNN-attention hybrid models that concatenate CNN with a self-attention mechanism have achieved cutting-edge results for numerous tasks such as acoustic event classification [17, 18] and other audio pattern recognition topics [19, 20]. Gong et al. [21] took one step further and devised purely attention-based models for audio classification. The devised Audio Spectrogram Transformer (AST) was assessed on various audio classification benchmarks with new state-of-the-art results, which shows that CNNs are not indispensable in this context.
@@ 2.1 Featurization
Similar to prior literature [11, 12], the Gammatone ERB filterbank is selected for generating time-frequency representation as it leads to low model complexity while preserving signal information that is relevant to this problem. Specifically, the Gammatone filterbank consists of 20 bands covering the range from 50Hz to 2000Hz. We compute the audio by continuing with a Hann window of 64 samples and a hop size of 32 samples, followed by convolution with the filterbank. This convolution generates a spectral feature (). Additionally, the phase information of the audio is retained. The phase angles computed for each time-frequency bin can be used to generate a phase feature, and it is then truncated to only include the bands corresponding to frequencies below 500 Hz (i.e. ) as lower frequency behavior generally carries more information corresponding to the room volume [6]. Furthermore, the first-order derivative of phase coefficients along the frequency axis is also concatenated (i.e. ). The configuration of this feature set aligns with the “+Phase” model outlined in [16], which is shown to outperform other methods that rely solely on magnitude-based spectral features.
@@ 4.1 Experimental Design
During the model training phase, MSE was used as the loss function and the Adam optimizer from PyTorch was deployed for optimization. The CNN model and the proposed method were trained for 1000 epochs and 150 epochs, respectively. This decision was made as good convergence behaviors were already observed for both training and validation at these epoch numbers. L2 regularization was applied to mitigate potential over-fitting. Additionally, an adaptive learning rate strategy was adopted to ensure efficient convergence during training. If the model failed to demonstrate improvement on the validation set for ten consecutive epochs, early stopping criteria were applied to halt the training process. For comparison test purposes, we switched between Dataset I and Dataset II, as well as whether or not to use the pretrained model from ImageNet, while maintaining same hyperparameters between the attention-based and CNN-based model to ensure uniformity in model configurations.
@@ 4.2.1 Base Systems
It can be seen that the proposed model significantly outperforms the CNN model in terms of prediction accuracy, relationship with ground truth values and predictive capability. This indicates that neural networks purely based on attention are sufficient (or even more superior than CNN models) to accurately learn the relationship between indoor acoustic characteristics and room volumes, even with the low-layer network configuration and a relatively small number of training epochs.
@@ 4.2.2 Enhanced Systems With Pretraining
With Dataset I, the deployment of the ImageNet pretraining elevated the proposed method’s performance to a new level, yielding a significantly improved room volume estimation accuracy. With the application of the SpecAugment method, Dataset II facilitated to further enhance system performance for all three models. Particularly, the augmentation effect was more prominent in the proposed method w/ Pretrain, confirming the effectiveness of SpecAugment in terms of alleviating overfitting and enhancing models’ generalizability. As a more illustrative example, the best-performing system, i.e. “Proposed method w/ Pretrain” model with Dataset II, resulted in a median and mean absolute error of only 155 and 1219 in linear scale respectively, given that the range of test set room volumes was [12, 21000] . In contrast, the median and mean absolute error of the CNN-based system with Dataset II was 353 and 1919 , respectively.